# Multi-perspective Alignment for Increasing Naturalness in Neural Machine Translation

### v1

@@ 6.4 Finer-grained Analysis
In Table 4, we compare the surface-level output of the strongest baseline (Tagging; 4.8M) with that of the base MT model and our alignment system. As highlighted in green, the English ‘community hikes’ is translated to gemeenschapsfietsen (‘community bicycles’) by the Tagging system, while our alignment system outputs gemeenschapshikes (‘community hikes’). This is an example of how the Tagging model output may score high on lexical diversity metrics, but strays from the content, where our model preserves it. As shown in blue, ‘general clean-mindedness’ is translated to algehele schoonheid (‘overall beauty’) by the base MT system. Our alignment system translates to algemene properheid (‘general cleanliness’), while the Tagging system outputs algemeene properheid. The latter case contains a double e, which is not typical in this context for modern Dutch, but does appear in the original Dutch dataset. Our alignment MT system is not affected by this.
@@
Figure 3 shows MTLD scores per book between human translation, base MT model, and alignment model (COMET + MT-HT). We observe that COMET + MT-HT scores are higher than the base MT model for all books, indicating that our alignment method makes the translations more lexically diverse. Interestingly, our method brings the results closer to or even exceeds HT in terms of lexical diversity on some books (e.g. 5, 9, 14, and 16). Overall, the MTLD scores of the alignment models are between those of the base MT model and human translation.
Figure 3: Per-book comparison of MTLD. Note that avg presents the average score across all books.
@@ 6.5 Impact of Hyper-parameter
To examine the impact of hyper-parameter $\beta$ (see Section 4.3), we report the results when it is set to 0.0, i.e. only considering the reward learning. Models trained without the constraint of negative log-likelihood loss lead, as expected, to worse content scores across the board as they move too far from the base MT model. On the other hand, these models achieve better classification scores but worse naturalness results (except B1 and CDU in MT-OR). The higher scores on classifiers could be due to characteristics of translated language beyond those related to high lexical diversity. Future work is needed to determine how the classifiers, lexical diversity, machine translationese and naturalness are precisely related.
@@ 7 Conclusion
We proposed a reinforcement learning based alignment framework for machine translation, which improves translation quality from multiple perspectives. Using the evaluation model COMET and different binary classifiers trained with MT, HT, and original target-language data as reward models, we approximate human preference and align the MT model with it. Our experiments on English-to-Dutch literary translation show that our model produces translations that are lexically richer and more natural without loss in translation accuracy.
8 Limitations
Due to the computational resources required to conduct this research, we were only able to perform extensive experiments on one language pair and domain. Since we first wanted to show that our method is sound in a simple setting, i.e. training a model from scratch, we have not proceeded to involve complex settings and computationally-heavy models, such as pre-trained large language models. Furthermore, our metrics for evaluating naturalness are mostly limited to lexical diversity, while writing style in general is much broader and difficult to capture with automatic metrics. We acknowledge that large-scale human evaluation, beyond our surface-level inspection in Section 6.4, could bring important insights.

### v2

@@ 6.4 Finer-grained Analysis
In Table 4, we compare the surface-level output of the strongest baseline (Tagging; 4.8M) with that of the base MT model and our best alignment system (COMET & MT-HT). As highlighted in green, the English ‘community hikes’ is translated to gemeenschapsfietsen (‘community bicycles’) by the Tagging system, while our alignment system outputs gemeenschapshikes (‘community hikes’). This is an example of how the Tagging model output may score high on lexical diversity metrics, but strays from the content, where our model preserves it. As shown in blue, ‘general clean-mindedness’ is translated to algehele schoonheid (‘overall beauty’) by the base MT system. Our alignment system translates to algemene properheid (‘general cleanliness’), while the Tagging system outputs algemeene properheid. The latter case contains a double e, which is not typical in this context for modern Dutch, but does appear in the original Dutch dataset. Our alignment MT system is not affected by this.
@@
Figure 4 shows MTLD scores per book between human translation, base MT model, and our best alignment model (COMET + MT-HT). We observe that COMET + MT-HT scores are higher than the base MT model for all books, indicating that our alignment method makes the translations more lexically diverse. It is interesting to see that our method brings the results closer to or even exceeds human translation in terms of lexical diversity on some books (e.g. 5, 9, 14, and 16). Overall, the MTLD scores of the alignment models are between those of the base MT model and human translation.
Figure 4: Per-book comparison of MTLD under human translation, base MT model, and alignment model. Note that avg presents the average score across all books.
@@ 6.5 Impact of Hyper-parameter
To examine the impact of hyper-parameter $\beta$ (see Section 4.3), we report the results when it is set to 0.0, i.e. only considering the reward learning. Models trained without the constraint of negative log-likelihood loss lead, as expected, to worse content scores across the board as they move too far from the base MT model (Table 5). These models achieve better classification scores but worse naturalness results (except MTLD, B1, and CDU in MT-OR and MTLD in MT-HT). The higher scores on classifiers could be due to characteristics of translated language beyond those related to high lexical diversity. Future work is needed to determine how the classifiers, lexical diversity, machine translationese and naturalness are precisely related.
@@ 7 Conclusion
We proposed a reinforcement learning based alignment framework for machine translation, which improves translation quality from multiple perspectives. Using the evaluation model COMET and different binary translationese classifiers trained with MT, HT, and original target-language data as reward models, we approximate human preference and align the MT model with it. Our experiments on English-to-Dutch literary translation show that our model produces translations that are lexically richer and more natural without loss in translation accuracy.
Limitations
Due to the computational resources required, we were only able to perform extensive experiments on one language pair and domain. Since we first wanted to show that our method is sound in a simple setting, i.e. training a model from scratch, we have not proceeded to involve complex settings and computationally-heavy models, such as pre-trained large language models. Furthermore, our metrics for evaluating naturalness are mostly limited to lexical diversity, while writing style in general is much broader and difficult to capture with automatic metrics. We acknowledge that large-scale human evaluation, beyond our surface-level inspection in Section 6.4, could bring important insights.
Acknowledgments
The anonymous reviewers of ACL Rolling Review provided us with useful comments which contributed to improving this paper and its presentation, so we’re grateful to them. We would also like to thank the Center for Information Technology of the University of Groningen for their support and for providing access to the high-performance computing cluster Hábrók. Huiyuan Lai is supported by the Sector Plan in the Humanities “Humane AI” funding from the Dutch Ministry. Antonio Toral is supported by a Beatriz Galindo senior fellowship (BG23/00152) from the Spanish Ministry of Science and Innovation.